# 0.5 eV QCD Axion Cosmology

### v48

@@ 1 Introduction
5 0.5 eV QCD Axion Structure Formation
Figure 1: Cooling of 0.5 eV QCD axion dark matter in the early universe shows a cross-over in the adiabatic compressibility, $dp_{A}/du_{A}$ (purple crosses), from the radiation-like value, $dp_{A}/du_{A}\approx 1/3$, at large redshift, to the matter-like value, $dp_{A}/du_{A}\approx 0$, at small redshift, starting right around the redshift of “matter-radiation equality,” $z_{\rm eq}=3~376~(19)$ (vertical orange arrow), predicted by the standard cold dark matter cosmology with cosmological constant ($\Lambda{\rm CDM}$) using observations of the early universe (Planck Collaboration et al., 2019, 2020).
The formation of large-scale astrophysical structures in the early universe is a key test of any proposal for the nature of dark matter, dark energy, and the Big Bang, including 0.5 eV QCD axion cyclic cosmology. Historically, the now-standard cold dark matter cosmology first emerged in the context of concerns about the relative lack of large-scale anisotropy in the cosmic microwave background (CMB), compared to expectations based on the then-standard baryon-based Big Bang cosmology (P. J. E. Peebles, 2020, pg. 301–302). In fact, Peebles first proposed a hot dark matter cosmology (P. J. E. Peebles, 1982a), before turning to cold dark matter (P. J. E. Peebles, 1982b), with the shared goal, in both cases, of “washing out” large-scale CMB anisotropy and, in this way, resolving the tension between the then-standard cosmology and early-universe observations by D. J. Fixsen et al. (1983).
As the 0.5 eV QCD axion dark matter equation of state crosses over to matter-like at small redshifts from radiation-like at large redshifts, starting around $z_{\rm eq}$, as shown in Fig. 2, the dynamics of the expansion change to “matter-dominated” from “radiation-dominated.” Indeed, with six axions for each photon made in the cosmic background radiation (See Section 3 and Appendix A), axions dominate the expansion dynamics on both sides of “matter-radiation equality.” The cross-over in the axion equation of state (Fig. 2), beginning around $z_{\rm eq}$, drives the cross-over in the character of the expansion which then drives the large-scale structure formation seen by early-universe observations, just as the change from “radiation-domination” to “matter-domination” drives such structure to form in the standard cold dark matter cosmology (P. J. E. Peebles, 1993, pgs. 624–625).
The rough agreement, shown in Fig. 2, between the best present observational value of the redshift where large-scale structure formation begins, $z_{\rm eq}$, and the predicted value for this redshift based on the cross-over in the equation of state for the 0.5 eV QCD axion dark matter, is a crucial test for the proposed cosmology. Together with the striking agreement between the predicted and observed values for the present energy density of cold dark matter in the universe (See Section 3), this agreement on large-scale structure formation provides compelling evidence that the proposed 0.5 eV QCD axion dark matter cosmology is compatible with early-universe observations. In this way, we deepen the proposed resolution of the Hubble tension, sketched in Section 1, by ensuring that the agreement between the standard $\Lambda{\rm CDM}$ cosmology and early-universe observations is not spoiled by the switch to “cooling” dark matter, in the form of 0.5 eV QCD axion dark matter with the equation of state shown in Fig. 2.
6 0.5 eV QCD Axion Nucleosynthesis
Figure 2: Cooling of the universe in the first three minutes after the Big Bang for a mix with six 0.5 eV QCD axions for each photon (curved line) and for the standard mix of photons and light leptons (filled and unfilled boxes) both reach the freeze-out of the neutron-to-proton ratio (horizontal line with dots) around the same time, but the standard mix (P. J. E. Peebles, 1966; G. Steigman, 2007) takes roughly twice as long to start forming helium nuclei (horizontal line with boxes).
Indeed, doubling the baryon number density, $n_{B}(T_{d})$, compared to the standard cold dark matter cosmology value, as we propose to do to resolve the Hubble tension (See Section 1 and 3), ensures a nearly complete conversion of deuterons into helium, and, thus, a nearly complete absence of primordial deuterium (R. J. Cooke et al., 2018), for the timing of the turn-off, $t_{d}(\rm{\Lambda CDM})\approx 200~{\rm sec}$ (P. J. E. Peebles, 1966; G. Steigman, 2007), set by the standard mix of light leptons and photons. Yet, with the appropriate mix of six axions for every photon made in the Big Bang (See Section 3 and Appendix A), the turn-off time, $t_{d}\approx 100~{\rm sec}$, comes nearly twice as fast as it does with the standard mix (See Fig. 3). Thus, the value of the product that determines the deuterium abundance, $n_{B}(T_{d})\langle\sigma v\rangle t_{d}\sim 1$, lands in about the same place, preserving the agreement with observations, where $\langle\sigma v\rangle$ is the thermal average of the product of the neutron capture cross-section, $\sigma$, with the nucleon relative velocity, $v$, (P. J. E. Peebles, 1993, pg. 141–142).
Still, the change in the time of the turn-off, $t_{d}$, has observable consequences due to the change in the gap in time, $t_{d}-t_{n}$, between the start of BBN, at $t_{d}$, and the “freeze-out,” at $t_{n}$, of the weak nuclear reactions between light leptons and nucleons (Fig. 3). During this gap, the decay of free neutrons with life-time, $\tau_{n}\approx 878~{\rm sec}$ (P. J. Mohr et al., 2025), cuts down the neutron-to-proton ratio (P. J. E. Peebles, 1993, pg. 184–186). By cutting $t_{d}$, while leaving $t_{n}$ more or less unchanged, we shorten the gap, $t_{d}-t_{n}$, by about ten percent of the free neutron lifetime, $\tau_{n}$, so that about ten per cent more neutrons make it to the start of BBN, at $t_{d}$, after which essentially all the surviving free neutrons end up as primordial helium (P. J. E. Peebles, 2020, pg. 128–129): We thus expect a roughly ten percent larger value for the primoridal helium abundance, $Y_{\rm P}\approx 0.27$, compared to the best present estimate with the standard mix of light leptons and photons, $Y_{\rm P}({\rm\Lambda CDM})\approx 0.25$ (C. Pitrou et al., 2018; S. Gariazzo et al., 2022).
7 0.5 eV QCD Axion Initial Fluctuations
More than five decades ago, E. R. Harrison (1970) and Y. B. Zeldovich (1972) argued for a scale-invariant initial spectrum of fluctuations in the energy density of the radiation created in the Big Bang, $P(k)=Ak^{n_{s}}$, where the spectral index value, $n_{s}=1$, ensures scale-invariance, and the constant, $A$, is independent of the co-moving wavevector of the fluctuation, $k$. Within the standard $\Lambda{\rm CDM}$ cosmology, the best present estimate from early-universe observations for the value of the spectral index, $n_{s}({\rm\Lambda CDM})=0.9677~(37)$ (Planck Collaboration et al., 2019, 2020), falls $8.8\sigma$ below the scale-invariant value, $n_{s}=1$. By contrast, the conformal symmetry of the cross-over from future infinity to the next big bang in 0.5 eV QCD axion cyclic cosmology (See Appendix C) requires the scale-invariant value, $n_{s}=1$, in agreement with the argument of E. R. Harrison (1970) and Y. B. Zeldovich (1972), but in sharp tension ($8.8\sigma$) with the standard $\Lambda{\rm CDM}$ value.
8 0.5 eV QCD Axion Astrophysical Constraints
Figure 3: Axion-photon coupling grows with initial core helium mass fraction in simulations of stellar evolution for horizontal branch stars constrained by observations of globular clusters (curved line with dots) (A. Ayala et al., 2014). The predicted value for the axion-photon coupling strength (horizontal lines with boxes), given in Eq. (7), combines with this empirical-computational relation to give a prediction for the helium mass fraction (box with whiskers) that differs by $5.0\sigma$ from the primordial helium abundance expected within the standard cold dark matter cosmology with cosmological constant ($\rm\Lambda CDM$}) based on simulations of Big Bang nucleosynthesis (thick vertical line) (S. Gariazzo et al., 2022; C. Pitrou et al., 2018), but that agrees rather well with the best present estimates for the initial core helium mass fraction in the Sun (pair of thin vertical lines) (L. Piersanti et al., 2007; A. M. Serenelli & S. Basu, 2010).
$R=6.26Y_{\rm HB}-0.41g_{10}^{2}-0.12,$ (32)
where $R=N_{\rm HB}/N_{\rm RGB}=1.39~(3)$ was the observed ratio of the number of HB stars, $N_{\rm HB}$, to the number of stars on the red giant branch (RGB) of the color-magnitude diagram, $N_{\rm RGB}$, in 39 globular clusters. Combining the empirical-computational relation in Eq. (32) with the predicted value for the axion-photon coupling in Eq. (7), we find a percent-level precise estimate for the initial core helium mass fraction in these HB stars that formed in the first billion years after the Big Bang:
$Y_{\rm HB}=0.272~(5)~[18~{\rm per~mille}],$ (33)
which differs by $5\sigma$ with the best present estimate of the “primordial” helium abundance, $Y_{P}({\rm\Lambda CDM})=0.247~09~(17)~[0.69~{\rm per~mille}]$ (C. Pitrou et al., 2018; S. Gariazzo et al., 2022), within the standard cold dark matter cosmology, but that roughly matches the best present estimates for the initial core helium mass fraction in the Sun, $Y_{\odot}=0.273~(6)~[{\rm 22~per~mille}]$ (A. M. Serenelli & S. Basu, 2010; L. Piersanti et al., 2007). As discussed in Section 6, the mix of axions and photons that dominates the expansion of the universe leading up to Big Bang nucleosynthesis, within the proposed physical picture, yields about ten percent more primordial helium, $Y_{P}\approx 0.27$, than the standard mix of light leptons and photons, and, in this way, resolves the $5\sigma$ tension between $Y_{\rm HB}$ in Eq. (33) and $Y_{P}({\rm\Lambda CDM})$.
Similarly, the recent attempts to “rule out” axion dark matter with the rest-mass energy in Eq. (6) and the axion-photon coupling in Eq. (7) using “blank-sky” observations of the dark matter “halo” of the Milky Way at the near-infrared wavelength $2\lambda_{A}=2h/m_{A}c\approx 4.92~{\rm\mu m}$, corresponding to the decay of the axion into a pair of photons, subtract a “smooth astrophysical background” by fitting a quartic polynomial to the near-infrared spectrum within a window of width $\Delta\lambda\approx 0.2~{\rm\mu m}$ (E. Pinetti, 2025). However, the proposed physical picture implies that 0.5 eV QCD axion dark matter is in thermal equilibrium with the cosmic microwave background (See Section 3 and Appendix A), and, thus, the axions in the dark matter “halo” of the Milky Way have velocity dispersion, $\sigma_{v}$, fixed by the CMB temperature, $T_{\gamma}$, and by the 0.5 eV QCD axion dark matter rest-mass energy, $m_{A}c^{2}$, leading to an expected line-width, $2\Delta\lambda_{A}$, that is roughly the same size as the window used to subtract the “smooth astrophysical background:”
$2\Delta\lambda_{A}$ $=$ $2\lambda_{A}\frac{\sigma_{v}}{c}$ (34)
$=$ $2\lambda_{A}\sqrt{\frac{3kT_{\gamma}}{m_{A}c^{2}}}$
$\approx$ $0.2~{\rm\mu m}.$
9 Conclusions
Among astrophysicists during the past forty years, three main “pillars” have come to be viewed as supporting the now-standard cold dark matter cosmology ($\rm\Lambda CDM$}): the anisotropy of the cosmic microwave background (CMB), the “primordial” abundance of chemical elements with low atomic mass whose nuclei were synthesized in the Big Bang (BBN), and the redshift-magnitude relation of Type Ia supernovae (P. J. E. Peebles, 2020, pg. 323–340). The physical picture of dark matter, dark energy, and the Big Bang, presented in this Letter, rests on the same “pillars” of support, but does so more naturally. In particular, the proposed physical picture naturally resolves the now more than $5\sigma$ tension in the present expansion rate of the universe between $\rm\Lambda CDM$} and “late” universe observations, and does so in favor of the observations.
Still, most of the cosmological parameters come out with around the same values as they do in $\rm\Lambda CDM$}. The energy density of cold dark matter in the present universe and the cosmological constant, for example, land within the uncertainties of the best present $\rm\Lambda CDM$} values from fits to early universe observations. However, it is crucial to emphasize that these essentially shared values arise as first-principles predictions, not fits to observations, within the proposed physical picture.
Similarly, the redshift of “matter-radiation equality” and the “primordial” deuterium abundance come out about the same as $\rm\Lambda CDM$}. This redshift is the critical “epoch” in the early universe where large-scale structure begins to form, and its value determines the location of the main “peak” in the angular power spectrum of the CMB fluctuations. Likewise, the primordial deuterium abundance is a key indicator of the timing and the conditions in the Big Bang when “light” element nucleosynthesis begins.
In conclusion, what the standard $\rm\Lambda CDM$} cosmology does well, the present physical picture does just as well or better. More sharply, the present physical picture provides an actual physical picture for the main ingredients of $\rm\Lambda CDM$}: dark matter, dark energy, and the Big Bang. In this sense, $\rm\Lambda CDM$} can be viewed, as it was intended to be viewed (P. J. E. Peebles, 1993, pg. 621–622), as a way of stimulating “the development of ideas on how to find and test models for the origin of [large-scale] structure,” rather than as an actual physical picture for the nature of dark matter (or anything else).
Moving forward, the sharp difference in the baryon abundance, roughly $190\sigma$, gives about a factor of two more baryons with the present physical picture. At the same time, the mix of six axions for every photon made in the Big Bang, within this picture, speeds up the cooling off of the universe in the first three minutes. The combination of these two sharp differences leads to about the same deuterium abundance but about $5\sigma$ more helium compared to $\rm\Lambda CDM$}, we have argued, and these loose arguments, in part based on astrophysical observations of globular clusters, must be confirmed by actual BBN calculations within the proposed physical picture.
In a similar vein, an actual calculation of the spectrum of CMB fluctuations within the proposed physical picture must be done. We have argued, indirectly, that the location of the main peak in the spectrum should come out about the same as it does in $\rm\Lambda CDM$} due to the coincidence of the redshift where the 0.5 eV QCD axion dark matter equation of state crosses over from radiation-like to matter-like, on the one hand, and, on the other, the redshift of “matter-radiation equality” within $\rm\Lambda CDM$}. Along similar lines, there is reason to believe that the sharp difference in baryon abundance, roughly $190\sigma$, combines with the sharp difference in the “primordial” spectral index, nearly $9\sigma$, to give basically the same spectrum as the best fit within $\rm\Lambda CDM$}, so that the agreement of this best fit with early universe observations is retained in the present physical picture, albeit with first-principles predictions taking the place of the key fit parameters.
Finally, the straightforward resolution of the Hubble tension using 0.5 eV QCD axion cyclic cosmology, presented in this Letter, suggests that we must “re-visit” many of the conventional assumptions about the way dark matter interacts with light energy and “bright” matter. For much of the past half century, it has been conventional to assume that dark matter “tracks” bright matter in gravitational equilibrium within a “halo” around the Milky Way galaxy, but the lack of a sharp two-photon decay “line” from 0.5 eV QCD axion dark matter in the infrared spectrum of this halo, suggests that dark matter is, instead, in thermal equilibrium with the photons of the CMB, as predicted by the proposed physical picture. For an equally long span of time, axions have been assumed to play essentially no role in the flow of energy within compact astrophysical bodies, such as pulsars, and within Type II supernovae, such as supernova 1987A, and, thus, it may be time to re-visit those assumptions, in light of the resolution of the Hubble tension presented in this Letter.
@@ 9 Conclusions
Appendix A 0.5 eV QCD Axion Self-Energy and Number Density
Within the physical picture presented in Section 2, the gravitational self-energy of 0.5 eV QCD axion dark matter, $U_{\Lambda}=N_{U}G(E/c^{2})^{2}/R$, given in Eq. (8), is generated by the kinetic energy, $E$, of a massless particle and anti-particle, each with de Broglie wavelength, $\lambda=hc/E=2\pi R$. The value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$, then follows from two assumptions. First, the axion is “spin-less,” in the sense that the plane of motion of the particle and anti-particle has equal probability amplitude to be oriented in any direction in three-dimensional space, and, second, the three-dimensional vector, $\bf r$}, separating the particle and anti-particle, is related to twice the de Broglie wavelength, $D=2\lambda$, by the familiar geometric relation, $1/D=\langle 1/|{\bf r}|\rangle_{\Omega}$, for the angle-averaged plane-projected (“line-of-sight”) distance (V. Barger & M. Olsson, 1995, pgs. 338–339):
@@
The predicted value for the rest-mass energy, $m_{A}c^{2}$, of the 0.5 eV QCD axion dark matter, given in Eq. (6), follows from the analysis of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, presented at the end of Section 3, once a relation can be found with the temperature of the Big Bang, $T_{\pi}$. Both of the quantities, $m_{A}c^{2}$ and $kT_{\pi}$, enter the statistical mechanical expression for this ratio, given in Eq. (24), through their ratios with the nucleon rest-mass energies, $m_{N}c^{2}$, for N=n,pN=n,~p. Given an independent relation between $m_{A}c^{2}$ and $kT_{\pi}$, then, the matching of this expression to the numerically exact value for the ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, given in Eq. (5) and derived in Section 4, determines both quantities with a precision that is limited only by that of this independent relation.
@@
Appendix C 0.5 eV QCD Axion Cyclic Cosmology
@@
where the value of the axion decay constant, $f_{A}\approx 10^{10}~{\rm MeV}$, given in Section 2, ensures the condition $\hat{\ell}\check{\ell}=1$, which has been suggested previously on geometric grounds (R. Penrose, 2016).
With this conformal interpretation, the cyclic cosmology matching condition, given in Eq. (29), simply expresses the conformal rescaling of the Ricci curvature scalar at the cross-over, $R$, to its value at future infinity, $\hat{R}$, and at the Big Bang, $\check{R}$, given by (R. Penrose & W. Rindler, 1986, pg. 123)
@@
where $u_{B}$ is the energy density of baryons in the present universe, $T_{\pi}$ is the Big Bang temperature, and $T_{\gamma}$ is the temperature of the cosmic microwave background. Combining Eq. (C), (C2), and (C), we arrive at the cyclic cosmology matching condition
@@
We estimate the small dimensionless quantity, $\epsilon=(4m_{\alpha}/m_{p}-1)Y_{P}=1.876~(34)\times 10^{-3}~[18~{\rm per~mille}]$, using the best present value for the ratio, $m_{\alpha}/m_{p}=3.972~599~690~252~(70)~[18~{\rm parts~per~trillion}]$ (P. J. Mohr et al., 2025), of the mass of the helium nucleus to the proton mass, and using the astrophysical determination of the primordial helium abundance within the present physical picture, $Y_{P}=Y_{\rm HB}=0.272~(5)~[18~{\rm per~mille}]$ (See Section 8). The exact expression for the relativistic correction, $1+\delta=m_{p}\eta/(m_{p}e^{-x_{p}}x_{p}J(x_{p})+m_{n}e^{-x_{n}}x_{n}J(x_{n}))$, uses the nucleon mass-to-temperature ratios at the Big Bang, $x_{N}=m_{N}c^{2}/kT_{\pi}$, for $N=p$ and $n$, with $kT_{\pi}$ from Eq. (3). The numerically exact integral, $J(x_{N})$, is defined in Eq. (25), while the integral, $I(x_{N})=(1/2!\zeta(3))\int_{0}^{\infty}dxe^{-x}(x+x_{N})\sqrt{x(x+2x_{N})}$, determines the baryon-to-photon number density ratio, $\eta=e^{-x_{p}}I(x_{p})+e^{-x_{n}}I(x_{n})=1.290~113~(5)\times 10^{-9}~[{\rm 4~ppm}].$
@@
The derivation of the expression for the cosmological parameter of the cosmological constant, $\Omega_{\Lambda}h^{2}$, in Eq. (27), begins with the physical picture for the nature of dark energy presented in Section 2, but evaluated now at future infinity instead of the Big Bang:
@@
where $\hat{U}_{\Lambda}=N_{U}\hat{E}^{3}/E_{\rm Pl}^{2}=N_{U}f_{A}$ is the dark energy inside each axion at future infinity, and $\hat{n}_{A}=(\check{\ell}/\hat{\ell})^{3}(T_{\pi}/T_{\gamma})^{3}n_{A}=N_{n}(\check{E}/\hbar c)^{3}(\check{E}/\hat{E})^{3}$ is the axion number density at future infinity. See Appendix A for the derivation of the dimensionless constants, $N_{U}=1/(2\pi^{2})$ and $N_{n}=12\zeta(3)/(\pi^{2}x_{k}^{3})$, but see Appendix C for the conformal scale factors, $\hat{\ell}=\hat{E}/f_{A}$ and $\check{\ell}=\check{E}/f_{A}$. The see-saw relation, $\check{E}\hat{E}=f_{A}^{2}$ (See Section 2), was used in the third line, while the relations, $\check{E}=x_{k}kT_{\pi}$, and, $m_{A}c^{2}f_{A}=\sqrt{\chi_{\rm QCD}}=(kT_{\pi}/y_{T})^{2}$, were used in the fourth line.

### v49

@@ I Introduction
IX Conclusions
We have presented a physical picture for the nature of dark matter, dark energy, and the Big Bang. Dark matter is made of quantum chromodynamic (QCD) axions with rest-mass energy around 0.5 eV. Dark energy is the gravitational attraction between the particle of matter and the particle of antimatter, both moving with the speed of light, that bind to make the dark matter in the moment of the Big Bang.
Before the Bang, massless quarks and leptons, together with their antiparticles, move in a critical soup, without either dark matter or dark energy. Correspondingly, at future infinity, the axion dark matter dissolves into this critical soup and the dark energy within the axion disappears as well. In between future infinity and the next big bang, the universe shrinks and the temperature rises until the baryon abundance at the next big bang and the cosmological constant satisfy the cosmological matching condition imposed by the conformal symmetry of the cross-over.
In the present universe, the rate with which space grows over time is predicted within the resulting 0.5 eV QCD axion cosmology with high-precision and from first-principles. Remarkably, the predicted value does not disagree sharply with the best present determination using late-universe observations of relatively nearby Type Ia supernovae with low redshift. In this sense, the proposed cosmology has the potential to resolve the Hubble tension between the now-standard cold dark matter cosmology and these late-universe observations.
However, to realize such a resolution, the proposed cosmology must also pass a battery of observational tests which draw on a wide range of astrophysical phenomena, including those at high redshift. Remarkably, the predicted values for the cosmological constant and matter abundance, from first-principles with high-precision, within 0.5 eV QCD axion cosmology, provide a pleasingly good description of the redshift-distance relation for high redshift Type Ia supernovae observations that stretch back into the middle-universe around the epoch of cosmic noon when quasars were at their most active and when stars were forming at their fastest rate. It is worth emphasizing that, while the predictions do indeed seem to fit the data quite well, the underlying cosmological parameters do not come from a fit to the data, but instead come from the physical picture for dark matter, dark energy, and the Big Bang presented here.
Along with the redshift-distance relation, the primordial deuterium abundance is a second key test of any proposed cosmology. Observations from cosmic noon of quasar absorption line spectroscopy provide a percent-level precise determination of this quantity. Again remarkably, the observed primordial deuterium abundance roughly matches the value expected within 0.5 eV QCD axion cosmology.
Similarly, we have argued that the proposed cosmology is consistent with estimates of the primordial helium abundance using observations of globular clusters and simulations of horizontal branch stellar evolution. The key point is that the core-cooling effects from 0.5 eV QCD axions must be included in the simulations. Doing so, one finds an estimate for the core helium mass fraction at formation for stars in globular clusters, which are among the oldest in the galaxy and which formed near cosmic noon, and this estimate roughly matches the value expected within 0.5 eV QCD axion cosmology for the primordial helium abundance.
Finally, early-universe observations provide a third crucial test for the proposed cosmology. Remarkably, again, the best early-universe observational estimate for the redshift when the expansion of the universe changes from radiation-dominated to matter-dominated seems to roughly agree with the value predicted within 0.5 eV QCD axion cosmology. Along with the agreement with observations from the late-universe and the middle-universe, the agreement with early-universe observations suggests that the proposed cosmology does in fact resolve the Hubble tension, and, moreover, it suggests that this 0.5 eV QCD axion cosmology offers substantial promise for a consistent and complete account of a wide range of cosmological observations throughout the life of the observable universe.
Concretely, a straightforward calculation of Big Bang nucleosynthesis within 0.5 eV QCD axion cosmology promises to yield high-precision predictions from first-principles for the primordial deuterium and helium abundances. Similarly, a straightforward calculation of large-scale structure formation within 0.5 eV QCD cosmology offers the prospect of estimating the spectrum of energy density fluctuations in the early-universe from first-principles with high-precision. In this way, the pleasing agreement, with late-universe observations on the Hubble constant and with middle-universe observations on the cosmological constant and the matter abundance, can be tested with observations that probe the early-universe astrophysical phenomena of Big Bang nucleosynthesis and large-scale structure formation.
Beyond these concrete and straightforward steps, which apply to any cosmology that proposes to resolve the Hubble tension, a more ambitious and targeted program of astrophysical activity comes to mind in light of the high-precision predictions from first-principles for the 0.5 eV rest-mass energy of the QCD axion dark matter and for the axion-photon coupling strength. Over the past half century, a wide range of astrophysical observations have been used to rule out, exclude, and constrain 0.5 eV QCD axion dark matter. We have argued that the analyses of these observations invoke astrophysical assumptions that must now be re-visited, in light of the remarkable agreement, summarized above, between 0.5 eV QCD axion cosmology and cosmological observations across the life of the observable universe.
For example, the core helium mass fraction at formation for stars that are now on the horizontal branch (HB) of globular cluster color-magnitude diagrams can be determined using stellar evolution simulations once one includes the core-cooling effect of the axion-photon coupling on the helium-burning core of HB stars. Using the high-precision prediction from first-principles for the value of the axion-photon coupling from 0.5 eV QCD axion cosmology, one finds a rough match between the initial core helium abundance of HB stars and the best present estimates of the initial core helium abundance of the Sun. This rough agreement suggests that the stellar core helium mass fraction at formation does not significantly change or evolve from cosmic noon to the late-universe, in contrast to the traditional astrophysical assumption that core helium abundance at formation grows with time.
To take a second case, the traditional astrophysical assumption is that dark matter sits in gravitational equilibrium with the bright tracers of galactic dynamics, such as OB associations, H II regions, and 21 cm clouds. Yet, observations of the blank-sky infrared spectrum of the Milky Way dark matter halo do not find the sharp line at 4.92 microns that would be expected from the spontaneous two-photon decay of 0.5 eV QCD axion dark matter under such an assumption. Instead, within 0.5 eV QCD axion cosmology, dark matter forms in the Big Bang in thermal equilibrium with the photons that now form the cosmic microwave background and this thermal equilibrium persists into the present universe.
A third case in point is the assumption that axions have an astrophysically significant coupling to electrons. Observations of the white dwarf luminosity function and the brightness of the tip of the red giant branch rule out this assumption within the physical picture presented here. Instead, as has been emphasized regularly over the past half century within the particle physics literature, axions have no direct coupling to electrons, and instead they couple to electrons only indirectly through the exchange of a pair of photons, suppressing the effects of axion-electron coupling below astrophysical uncertainties.
But perhaps the most fundamental such astrophysical assumption regarding axions in need of re-visitation, however, is the largely unstated one that any proposed astrophysical effect of 0.5 eV QCD axion dark matter must be sub-dominant to effects relying solely on particles from within the standard model of particle physics. When one sets aside this prejudice, a wide range of long-standing tensions, prominent crises, and puzzling phenomena present themselves as potential probes of the astrophysical effects of dark matter within the proposed cosmology. Although less concrete and more open-ended, the resulting program of astrophysical activity promises to bring wide-ranging progress and insight.
On this point, x-ray observations of pulsars and the neutrino flux detection from the Type II supernova 1987A (SN 1987A) are germane. Quantum chromodynamic axions with rest-mass energy around 0.5 eV couple strongly to nucleons: The predicted axion-nucleon coupling strength is roughly one thousand times stronger than the axion-photon coupling, for example. The fact that pulsars are still x-ray bright long after they form and the fact that neutrinos were observed at all from SN 1987A rule out a sub-dominant role for 0.5 eV QCD axions in the flow of energy within pulsars and Type II supernovae like SN 1987A.
Instead, within the physical picture presented here, axions with rest-mass energy 0.5 eV and with the corresponding axion-nucleon coupling predicted by QCD must dominate the dynamics of compact astrophysical objects, such as pulsars and the collapsing stellar core that leads to a Type II supernova such as SN 1987A. As pointed out by Turner right after the neutrino flux detection from SN 1987A, dense nuclear material is axionically thick, in the same sense that dense atomic material is optically thick. In this thick limit, the emission of energy from the compact astrophysical object becomes dominated by an axi-sphere that is analogous to the photosphere at the surface of stars.
In conclusion, the physical picture for the nature of dark matter, dark energy, and the Big Bang, presented in this Letter, raises important questions across a wide-range of astrophysical phenomena. These astrophysical questions are rooted in the apparent success of 0.5 eV QCD axion cosmology in resolving the Hubble tension. Still, concrete and straightforward calculations of Big Bang nucleosynthesis and large-scale structure formation must now be done, alongside the more open-ended and ambitious astrophysical activity suggested by this physical picture, to solidify the proposed resolution of one of the most important and pressing problems in cosmology.
@@ IX Conclusions
Appendix A 0.5 eV QCD Axion Dark Energy and Number Density
Within the physical picture presented in Section VI, the gravitational dark energy of 0.5 eV QCD axion dark matter, $U_{\Lambda}=N_{U}G(E/c^{2})^{2}/R$, given in Eq. (14), is generated by the kinetic energy, $E$, of a massless particle and anti-particle, each with de Broglie wavelength, $\lambda=hc/E=2\pi R$. The value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$, then follows from two assumptions. First, the axion is “spin-less,” in the sense that the plane of motion of the particle and anti-particle has equal probability amplitude to be oriented in any direction in three-dimensional space, and, second, the three-dimensional vector, $\bf r$}, separating the particle and anti-particle, is related to twice the de Broglie wavelength, $D=2\lambda$, by the familiar geometric relation, $1/D=\langle 1/|{\bf r}|\rangle_{\Omega}$, for the angle-averaged plane-projected (“line-of-sight”) distance (V. Barger & M. Olsson, 1995, pgs. 338–339):
@@
The predicted value for the rest-mass energy, $m_{A}c^{2}$, of the 0.5 eV QCD axion dark matter, given in Eq. (9), follows from the analysis of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, presented at the end of Section VII, once a relation can be found with the temperature of the Big Bang, $T_{\pi}$. Both of the quantities, $m_{A}c^{2}$ and $kT_{\pi}$, enter the statistical mechanical expression for this ratio, given in Eq. (30), through their ratios with the nucleon rest-mass energies, $m_{N}c^{2}$, for N=n,pN=n,~p. Given an independent relation between $m_{A}c^{2}$ and $kT_{\pi}$, then, the matching of this expression to the numerically exact value for the ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, given in Eq. (7) and derived in Section VIII, determines both quantities with a precision that is limited only by that of this independent relation.
@@
Appendix C 0.5 eV QCD Axion Cosmology
@@
where the value of the axion decay constant, $f_{A}\approx 10^{10}~{\rm MeV}$, given in Section VI, ensures the condition $\hat{\ell}\check{\ell}=1$, which has been suggested previously on geometric grounds (R. Penrose, 2016).
With this conformal interpretation, the cosmology matching condition, given in Eq. (35), simply expresses the conformal rescaling of the Ricci curvature scalar at the cross-over, $R$, to its value at future infinity, $\hat{R}$, and at the Big Bang, $\check{R}$, given by (R. Penrose & W. Rindler, 1986, pg. 123)
@@
where $u_{B}$ is the energy density of baryons in the present universe, $T_{\pi}$ is the Big Bang temperature, and $T_{\gamma}$ is the temperature of the cosmic microwave background. Combining Eq. (C), (C2), and (C), we arrive at the cosmology matching condition
@@
We estimate the small dimensionless quantity, $\epsilon=(4m_{\alpha}/m_{p}-1)Y_{P}=1.876~(34)\times 10^{-3}~[18~{\rm per~mille}]$, using the best present value for the ratio, $m_{\alpha}/m_{p}=3.972~599~690~252~(70)~[18~{\rm parts~per~trillion}]$ (P. J. Mohr et al., 2025), of the mass of the helium nucleus to the proton mass, and using the astrophysical determination of the primordial helium abundance within the present physical picture, $Y_{P}=Y_{\rm HB}=0.272~(5)~[18~{\rm per~mille}]$ (See Section V). The exact expression for the relativistic correction, $1+\delta=m_{p}\eta/(m_{p}e^{-x_{p}}x_{p}J(x_{p})+m_{n}e^{-x_{n}}x_{n}J(x_{n}))$, uses the nucleon mass-to-temperature ratios at the Big Bang, $x_{N}=m_{N}c^{2}/kT_{\pi}$, for $N=p$ and $n$, with $kT_{\pi}$ from Eq. (VII). The numerically exact integral, $J(x_{N})$, is defined in Eq. (31), while the integral, $I(x_{N})=(1/2!\zeta(3))\int_{0}^{\infty}dxe^{-x}(x+x_{N})\sqrt{x(x+2x_{N})}$, determines the baryon-to-photon number density ratio, $\eta=e^{-x_{p}}I(x_{p})+e^{-x_{n}}I(x_{n})=1.290~113~(5)\times 10^{-9}~[{\rm 4~ppm}].$
@@
The derivation of the expression for the cosmological parameter of the cosmological constant, $\Omega_{\Lambda}h^{2}$, in Eq. (33), begins with the physical picture for the nature of dark energy presented in Section VI, but evaluated now at future infinity instead of the Big Bang:
@@
where $\hat{U}_{\Lambda}=N_{U}\hat{E}^{3}/E_{\rm Pl}^{2}=N_{U}f_{A}$ is the dark energy inside each axion at future infinity, and $\hat{n}_{A}=(\check{\ell}/\hat{\ell})^{3}(T_{\pi}/T_{\gamma})^{3}n_{A}=N_{n}(\check{E}/\hbar c)^{3}(\check{E}/\hat{E})^{3}$ is the axion number density at future infinity. See Appendix A for the derivation of the dimensionless constants, $N_{U}=1/(2\pi^{2})$ and $N_{n}=12\zeta(3)/(\pi^{2}x_{k}^{3})$, but see Appendix C for the conformal scale factors, $\hat{\ell}=\hat{E}/f_{A}$ and $\check{\ell}=\check{E}/f_{A}$. The see-saw relation, $\check{E}\hat{E}=f_{A}^{2}$ (See Section VI), was used in the third line, while the relations, $\check{E}=x_{k}kT_{\pi}$, and, $m_{A}c^{2}f_{A}=\sqrt{\chi_{\rm QCD}}=(kT_{\pi}/y_{T})^{2}$, were used in the fourth line.